OpenStudy (anonymous):

I DON'T UNDERSTAND WHY MY ANSWER IS NOT CORRECT :( SLOPE IS -9/11 AND THE EQUATION IS -9X/11+44 Find the equation of the line parallel to -9x-11y=-2 and passing through the point (5,-1). The equation of the line is ??

OpenStudy (mertsj):

The slope of -9x-11y=-2 is -9/11 So plug that slope and the given point into point-slope form: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

i did..

OpenStudy (whpalmer4):

@mathcalculus you did your algebra incorrectly, it appears...

OpenStudy (anonymous):

can you point it out where?

OpenStudy (mertsj):

\[y-(-1)=\frac{-9}{11}(x-5)\]

OpenStudy (mertsj):

Simplify and you will be done.

OpenStudy (anonymous):

i have this: y=9x/11 +44

OpenStudy (whpalmer4):

your y-intercept value will be a fraction...maybe you didn't multiply the -9/11 when expanding

OpenStudy (anonymous):

@whpalmer4 i'm not sure where I have made my mistake.

OpenStudy (whpalmer4):

\[1+y = \frac{1}{11}(45-9x)\]probably screwed up bringing the 1 across

OpenStudy (mertsj):

\[y+1=\frac{-9}{11}x+\frac{45}{11}\]

OpenStudy (anonymous):

isn't y- (-1) a y+1?

OpenStudy (anonymous):

so if i bring it to the other side it must be subtracted from the 45

OpenStudy (anonymous):

oh yes! got it

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

9/11 * 5

OpenStudy (mertsj):

\[y=\frac{-9}{11}x+\frac{45}{11}-1=\frac{-9}{11}x+\frac{34}{11}\]

OpenStudy (whpalmer4):

subtract 11/11 from each side: \[y+1 - 11/11 = \frac{1}{11}{45-11-9x}\]\[y = \frac{1}{11}(34-9x)\]\[y = -\frac{9x}{11}+\frac{34}{11}\]

OpenStudy (whpalmer4):

whoops, used invisible () in that first eq. :-)

OpenStudy (anonymous):

that's not the answer..... :(

OpenStudy (anonymous):

it keeps saying it's wrong.

OpenStudy (anonymous):

i think it's the website, it must be correct. thank you everyone.

OpenStudy (whpalmer4):

well, that in fact is the equation of a line parallel to the original line, passing through -5,1 as you can see in this graph:

OpenStudy (mertsj):

Check the rules for entry. Maybe you are entering the fractions incorrectly or maybe you are not supposed to write y = or maybe they want the equation in some other form besides slope-intercept form.

OpenStudy (whpalmer4):

(5,-1) that is

OpenStudy (anonymous):

yes, I checked. appreciate the help

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

it must be correct, even the calculator shows it.

OpenStudy (mertsj):

Maybe you want to try 9x+11y=34

OpenStudy (whpalmer4):

I'll sometimes multiply by the denominator of the fraction, do the problem without the pesky fractions and accompanying mistakes, then divide again by the former denominator for the final answer.

OpenStudy (whpalmer4):

I'd say you know what you're doing, in any case...watch for those fraction mistakes :-)

OpenStudy (anonymous):

thanks alot ! :)

OpenStudy (mertsj):

yw